OpenStudy (anonymous):

Fe203 + 3CO produces 2Fe + 3C02 a 20.0 sample of iron 3 oxide reacts with 16.0 L of carbon monoxide at 1.50 atim at 200 c. The reaction produces 11.5 g of iron. Determine the percent yeild

OpenStudy (anonymous):

Fe2O3 + 3CO -> 2Fe + 3CO2 first we have to find moles(amount of substances) of Fe2O3 and CO, to do that: n(Fe2O3) = 20g/160g/mol = 0.125 mole to find mole of CO we use:PV=nRT(T = 200+273 = 473K) 1.5atm*16L = n*0.082*473K n = 0.62 mole of CO Fe2O3 is limiting..so n(Fe) = 2*0.125 = 0.25 mole of Fe m(Fe) = 0.25 * 56 = 14 g percent yield = 11.5/14 * 100 ~ 82 %

OpenStudy (anonymous):

can you help me with another problem